solve for x in simplest a+bi form:4x^2+17=4x
Answer by Edwin McCravy(20086)


solve for x in simplest a+bi form:4x^2+17=4x


                   4x² + 17 = 4x 

Get 0 on the right by adding -4x to both sides


              4x² - 4x + 17 = 0

Use the quadratic formula:
                  ______ 
            -b ± Öb²-4ac
        x = —————————————
                2a 

where a = 4; b = -4; c = 17

                      ______________ 
             -(-4) ± Ö(-4)²-4(4)(17)
        x = —————————————————————————
                     2(4) 
                  _______ 
             4 ± Ö16-272
        x = ——————————————
                  8

                  ____ 
             4 ± Ö-256
        x = ———————————
                 8

                   ___ 
             4 ± iÖ256
        x = ——————————
                8 

                   
             4 ± i(16)
        x = —————————
                8

            
             4 ± 16i
        x = —————————
                8

         
             4     16i     
        x = ——— ± —————
             8      8

             1     2i     
        x = ——— ± —————
             2      1

             1          
        x = ——— ± 2i
             2             


Edwin